OpenStudy (anonymous):

Which is an equation of a circle with center (-5, -7) that passes through the point (0, 0)? will give medal to who helps @perl

OpenStudy (perl):

The equation of a circle is : $$ \Large (x-h)^2 + (y-k)^2 = r^2 $$

OpenStudy (perl):

the center is (h,k)

OpenStudy (star109):

To write the equation of a circle, you have to have 2 things. You have to have the center of the circle and the radius. In the problem they give you the center of the circle. Now you need the radius of the circle. To find the radius of the circle you can use the distance formula. Once you use the distance formula you will find that the distance between the two points (0,0) and (-5,-7) is \[\sqrt{74}\]. Now you can write the equation of the circle. \[(x--5)^{2} + (y--7)^{2}= (\sqrt{74)^{2}}\] Now, let's simplify the equation (clean it up) \[(x+5)^{2}+ (y+7)^{2} = 74\]